OpenStudy (vera_ewing):

A solution has [OH-] = 3.5 × 10-6. Based on that, what must be true about this solution? A. It is a basic solution. B. It is an acidic solution. C. It is a strong acid. D. It is a strong base.

OpenStudy (vera_ewing):

@Australopithecus It's a solution right?

OpenStudy (australopithecus):

You can use, pOH = log[-OH] pH = log[+H] (note [H3O+] = [H+]) And 14 = pH + pOH pH scale here: https://encrypted-tbn3.gstatic.com/images?q=tbn:ANd9GcRItdnE1nXkLZef0p_sBk99e563UyVvPpltQusWPmBbI1cNELC-ZVkMazc

OpenStudy (australopithecus):

http://www.edu.pe.ca/gulfshore/Archives/ACIDSBAS/phscale2.jpg sorry use this link instead other is just a thumbnail

OpenStudy (vera_ewing):

So it's a strong base?

OpenStudy (australopithecus):

Note that the that is log base 10 not ln

OpenStudy (vera_ewing):

Ohh so strong acid?

OpenStudy (australopithecus):

yeah http://www.wolframalpha.com/input/?i=x+%2B+%28-log%283.5*10^-6%29%29+%3D+14 where x is pH

OpenStudy (vera_ewing):

It was A.

OpenStudy (vera_ewing):

It is a basic solution. Calculate the pOH using pOH = -log [OH-] = -log (3.5 × 10-6) = 5.5. This corresponds to a pH of 8.5. Because it is greater than 7, it is basic.

OpenStudy (australopithecus):

sorry wolfram alpha uses natural log by default

OpenStudy (australopithecus):

I told you it was log base 10 though

OpenStudy (australopithecus):

sorry for that

OpenStudy (vera_ewing):

Don't worry about it. I'm glad you helped.